# Learning Exhaustive Correlation for Spectral Super-Resolution: Where Unified Spatial-Spectral Attention Meets Mutual Linear Dependence

### 20 Dec 2023 v1

@@ Abstract
Spectral super-resolution from the easily obtainable RGB image to hyperspectral image (HSI) has drawn increasing interest in the field of computational photography. The crucial aspect of spectral super-resolution lies in exploiting the correlation within HSIs. However, two types of bottlenecks in existing Transformers limit performance improvement and practical applications. First, existing Transformers often separately emphasize either spatial-wise or spectral-wise correlation, disrupting the 3D features of HSI and hindering the exploitation of unified spatial-spectral correlation. Second, the existing self-attention mechanism learns the correlation between pairs of tokens and captures the full-rank correlation matrix, leading to its inability to establish mutual linear dependence among multiple tokens. To address these issues, we propose a novel Exhaustive Correlation Transformer (ECT) for spectral super-resolution. First, we propose a Spectral-wise Discontinuous 3D (SD3D) splitting strategy, which models unified spatial-spectral correlation by simultaneously utilizing spatial-wise continuous splitting and spectral-wise discontinuous splitting. Second, we propose a Dynamic Low-Rank Mapping (DLRM) model, which captures mutual linear dependence among multiple tokens through a dynamically calculated low-rank dependence map. By integrating unified spatial-spectral attention with mutual linear dependence, our ECT can establish exhaustive correlation within HSI. The experimental results on both simulated and real data indicate that our method achieves state-of-the-art performance. Codes and pretrained models will be available later.
1 Introduction
@@
Hyperspectral image (HSI) consists of multiple channels, with each channel representing the response in a specific spectral band. In comparison to the 3-channel RGB image, HSI excels in capturing detailed spectral information from a scene. Owing to this advantage, HSI finds extensive applications in image classification [12, 23, 36], object detection [30], face recognition [57], and more. However, acquiring 3D HSI with 2D sensors is challenging due to the mismatch of dimensions. Traditional scanning-based methods typically require multiple exposures to capture a full HSI, which is disadvantageous for dynamic and rapidly changing scenes.
@@
To address this issue, researchers have designed snapshot compressive imaging (SCI) systems with customized optical modulation and reconstruction algorithms, enabling snapshot acquisition of HSI [52, 47, 43, 8, 26]. However, these methods are often expensive and bulky in system implementation. Consequently, the task of HSI reconstruction from the easily obtainable RGB image [54], known as spectral super-resolution, has emerged as a popular solution.
@@
The crucial aspect of spectral super-resolution lies in exploiting correlations within HSI. Early research utilizes sparse coding [3] or low-rank representation [48] for spectral super-resolution. However, these methods often suffer from limited expressive power and generalization ability, thus failing to achieve satisfactory results. With the increasing computing power, learning-based methods have made significant progress in recent years and have become the mainstream solution for spectral super-resolution. Currently, Transformers [9, 40] have attained the state-of-the-art performance for spectral super-resolution by leveraging spectral-wise correlation through a spectral-wise self-attention mechanism. However, two types of bottlenecks exist that limit performance improvement and practical applications. First, existing Transformers predominantly focus on spectral-wise correlation while overlooking spatial-wise correlation in spectral super-resolution. Some works in other tasks [27, 15, 41, 32] attempt to model both spectral-wise and spatial-wise correlations together but often utilize separate network modules. The neglect and separation undermine the 3D nature of HSI and hinder the exploitation of unified spectral-spatial correlation. Second, the existing spectral-wise self-attention mechanism always learns the correlation between pairs of spectral bands, i.e. tokens, and captures the full-rank correlation matrix in the Transformer. These characteristics result in the inability to establish mutual linear dependence among multiple tokens.
@@
In this paper, we propose a novel Exhaustive Correlation Transformer (ECT) to model the unified spatial-spectral correlation and mutual linear dependence, both of which we believe are crucial for spectral super-resolution. The first motivation behind our method stems from the spatial-spectral similarity in HSI. Thus, we propose a Spectral-wise Discontinuous 3D (SD3D) splitting strategy to simultaneously model unified attention along the spectral and spatial dimensions. The SD3D splitting strategy contains continuous splitting in the spatial dimension and discontinuous splitting in the spectral dimension, allowing for an effective focus on spectral-wise non-local features without disrupting the continuous structure in the spatial dimension. The second motivation behind our methods arises from the information redundancy in HSI. Thus, we propose a Dynamic Low-Rank Mapping (DLRM) module to capture the mutual linear dependence among multiple tokens. The DLRM module simultaneously interacts among multiple tokens and maps them into a low-rank space, thereby learning a low-rank dependence map. By integrating unified attention with mutual linear dependence, our ECT can establish the exhaustive correlation within HSI and achieves state-of-the-art in extensive experiments on simulated and real data.
@@
[leftmargin=0.7cm]
We propose a novel Exhaustive Correlation Transformer (ECT) to model unified spatial-spectral correlation and mutual linear dependence for spectral super-resolution.
@@
We propose a Dynamic Low-Rank Mapping (DLRM) module to model the mutual linear dependence within HSI by dynamically calculating a low-rank dependence map among multiple tokens.
The experimental results indicate that our method achieves state-of-the-art performance on both simulated and real data, with the lowest error achieved under the lowest computational costs and the smallest number of parameters.
@@ 2.1 Spectral Reconstruction
HSI acquisition is typically carried out using push-broom cameras, which is time-consuming and challenging to capture dynamic or rapidly changing scenes. To address this issue, coded aperture snapshot spectral imaging (CASSI) systems have been widely used, generating 2D measurements [52, 34], which are then processed through a series of reconstruction algorithms [47, 42, 43, 8, 7, 10, 26, 20, 11, 51, 50] to obtain HSI.
However, CASSI systems are often expensive. Reconstructing HSIs from RGB images is a cost-effective alternative. Arad et al. [3] employ sparse coding for spectral super-resolution, while Aeschbacher et al. [1] use shallow learning models and achieve improved results. Due to the presence of substantial redundant information in HSI, a low-rank prior is critical for spectral reconstruction. There are several spectral reconstruction works [56, 55, 22, 48, 17, 16, 18] inspired by the low-rank prior. Recently, Three spectral super-resolution challenges [3, 4, 5] are held and significantly inspire the research. With the development of deep learning, convolutional neural networks are widely used in the spectral super-resolution task. Shi et al. [37] propose a convolutional neural network for spectral super-resolution, which win the NTIRE 2018 Challenge on Spectral Reconstruction from RGB Images. Li et al. [24, 25] introduce the channel attention mechanism into the convolutional neural network to improve the performance. Thanks to dynamic weights and long-range correlation modeling, Cai et al. [9] are the first to introduce Transformers into the field of spectral super-resolution and win first place in the NTIRE 2022 challenge [5].
@@ 3 Method
In this section, we first introduce the macro design of the Exhaustive Correlation Transformer (ECT). Then, we delve into the micro design within ECT.
@@ 3.2 Micro Design
As depicted in Figure 3, an Exhaustive Self-Attention Block (ESAB) comprises Exhaustive Self-Attention (ESA), Layer Normalization, and a Feed-Forward Network. The key process in ESA is summarized as follows: Firstly, the Spectral-wise Discontinuous 3D (SD3D) splitting strategy is applied to generate tokens, facilitating the exploitation of unified spatial-spectral correlation. Following that, the Low-Rank Dependence Map is generated through the Dynamic Low-Rank Mapping (DLRM) module to model mutual linear dependence among multiple tokens. The Full-Rank Attention Map is generated through the Unified Spatial-Spectral self-Attention (USSA) module to model the independent correlation between pairs of tokens. Next, we introduce the implementation details of ESA.
@@
where denotes softmax function and is a learnable parameter. , , and . normalization is performed within each Token to stabilize the training, and then the expressive power is improved by the learnable parameter . Since the dimension of tokens is greater than the number of tokens in this scenario and the Softmax operation after the dot product, there is a lower risk of rank reduction in the attention maps, as discussed in [6]. Typically, the learned attention maps have a full rank or nearly full rank. From the optimization point of view, since the linear correlation is different in different HSIs, the loss can only be minimized when the attention mechanism learns full-rank or nearly full-rank attention maps. The experiments confirm this point as well, and we will discuss it in the supplementary material. Furthermore, the scaled dot-product attention is calculated independently between paired tokens and cannot capture the mutual linear dependence among multiple tokens.
To address the limitation of self-attention in modeling mutual linear dependence within HSI, we propose a Dynamic Low-Rank Mapping (DLRM) module. The details of DLRM are illustrated in Figure 4. Initially, the tokens from the multi-head () are restored in a 3D manner and subsequently spatially pooled, reducing the spatial dimensions of the tokens from to . Following this step, the tokens are flattened in a 3D fashion to form a two-dimensional matrix. Finally, interactions take place among various heads and tokens through a 1d convolution, yielding a feature () with dimensions , where is a hyperparameter and . The matrices and then undergo a Softmax function, followed by transposition and multiplication to generate a dynamic matrix, which is a low-rank matrix with a rank no greater than . The calculation of the low-rank dependence map in DLRM is expressed by
The difference between the Attention Map in self-attention and the Dependence Map in DLRM is illustrated in Figure 5. As shown in the figure, the calculation of the correlation in self-attention is token-to-token, with each row of the correlation matrix independently calculated. Each element in the matrix is solely associated with two tokens. In contrast, DLRM first facilitates information exchange among various tokens and attention heads. Therefore, each element in the Dependence Map gathers information from multiple tokens. Consequently, each element in the Dependence Map aggregates information from multiple tokens. Moreover, the Dependence Map can implicitly model the linear correlations among multiple tokens due to the low-rank nature. In summary, the Dependence Map can effectively capture the mutual linear dependence among multiple tokens, which is absent in the Attention Map.
@@
| Method | NTIRE 2022 [5] | NTIRE 2020 [4] | ICVL [3] | Params | FLOPs | |||
| MRAE | RMSE | MRAE | RMSE | MRAE | RMSE | |||
| HSCNN+ [37] | 0.3814 | 0.0588 | 0.0684 | 0.0182 | 0.2322 | 0.0424 | 4.65 M | 304.45 G |
| HR Net [59] | 0.3476 | 0.0550 | 0.0682 | 0.0179 | 0.1139 | 0.0313 | 31.70 M | 163.81 G |
| AWAN [24] | 0.2191 | 0.0349 | 0.0668 | 0.0175 | 0.1040 | 0.0252 | 4.04 M | 270.61 G |
| Restormer [53] | 0.1833 | 0.0274 | 0.0645 | 0.0157 | 0.0945 | 0.0230 | 15.11 M | 93.77 G |
| DRCR [25] | 0.1823 | 0.0288 | 0.0664 | 0.0171 | 0.0763 | 0.0164 | 9.48 M | 586.61 G |
| MST-L [8] | 0.1772 | 0.0256 | 0.0646 | 0.0162 | 0.0728 | 0.0155 | 2.45 M | 31.82 G |
| MST++ [9] | 0.1645 | 0.0248 | 0.0624 | 0.0155 | 0.0691 | 0.0144 | 1.62 M | 22.29 G |
| ECT (Ours) | 0.1564 | 0.0236 | 0.0595 | 0.0146 | 0.0635 | 0.0142 | 1.19 M | 16.75 G |
@@ 4.1 Dataset
For the real data experiments, we capture real RGB images using FLIR Blackfly S BFS-U3-31S4 and obtained the spectral curves of the regions for validation using Specim IQ. We use HSI from the NTIRE 2022 training set and simulate RGB images to create paired training data. RGB simulation and more details about the datasets can be found in the supplementary material.
@@ 4.2 Implementation Details
For the evaluation metrics, following the NTIRE challenge, we use the Mean Relative Absolute Error (MRAE) and Root Mean Squared Error (RMSE) metrics to evaluate the performance. We primarily use MRAE as the main metric and also adopt it as the training objective. RMSE is used as an auxiliary metric.
For the network training details, we utilize a batch size of 40 and employ a learning rate schedule that follows the cosine annealing scheme, decreasing from 4e4 to 1e6 over 300 iterations. We choose the AdamW optimizer with parameters , , and weight decay is set to 1e4. RGB images are first split into patches and undergo random rotations and flips for data augmentation before being input into the network. The codes and pretrained models will be made publicly available.

### 18 Mar 2024 v2

# Learning Exhaustive Correlation for Spectral Super-Resolution: Where Spatial-Spectral Attention Meets Linear Dependence

@@ Abstract
Spectral super-resolution that aims to recover hyperspectral image (HSI) from easily obtainable RGB image has drawn increasing interest in the field of computational photography. The crucial aspect of spectral super-resolution lies in exploiting the correlation within HSIs. However, two types of bottlenecks in existing Transformers limit performance improvement and practical applications. First, existing Transformers often separately emphasize either spatial-wise or spectral-wise correlation, disrupting the 3D features of HSI and hindering the exploitation of unified spatial-spectral correlation. Second, existing self-attention mechanism always establishes full-rank correlation matrix by learning the correlation between pairs of tokens, leading to its inability to describe linear dependence widely existing in HSI among multiple tokens. To address these issues, we propose a novel Exhaustive Correlation Transformer (ECT) for spectral super-resolution. First, we propose a Spectral-wise Discontinuous 3D (SD3D) splitting strategy, which models unified spatial-spectral correlation by integrating spatial-wise continuous splitting strategy and spectral-wise discontinuous splitting strategy. Second, we propose a Dynamic Low-Rank Mapping (DLRM) model, which captures linear dependence among multiple tokens through a dynamically calculated low-rank dependence map. By integrating unified spatial-spectral attention and linear dependence, our ECT can model exhaustive correlation within HSI. The experimental results on both simulated and real data indicate that our method achieves state-of-the-art performance. Codes and pretrained models will be available later.
Keywords:
Spectral super-resolution Correlation Transformer1 Introduction
@@
Hyperspectral image (HSI) consists of multiple channels, with each channel representing the response in a specific spectral band. In comparison to the 3-channel RGB image, HSI excels in capturing detailed spectral information from a scene. Owing to this advantage, HSI finds extensive applications in image classification [12, 24, 37], object detection [31], face recognition [57], and more.
However, acquiring 3D HSI with 2D sensors is challenging due to the mismatch of dimensions. Traditional scanning-based methods typically require multiple exposures to capture a full HSI, which is disadvantageous for dynamic and rapidly changing scenes.
@@
To address this issue, researchers have designed snapshot compressive imaging (SCI) systems with customized optical modulation and reconstruction algorithms, enabling snapshot acquisition of HSI [51, 48, 44, 8, 27]. However, these methods are often expensive and bulky in system implementation. Consequently, the task of HSI reconstruction from the easily obtainable RGB image [54, 23], known as spectral super-resolution, has emerged as a popular solution with the advantages of being cheap and lightweight.
The crucial aspect of spectral super-resolution lies in exploiting correlations within HSI. Early research utilizes sparse coding [3] or low-rank representation [49] for spectral super-resolution. However, these methods often suffer from limited expressive power and generalization ability, thus failing to achieve satisfactory results.
@@
With the increasing computing power, learning-based methods have made significant progress in recent years and have become the mainstream solution for spectral super-resolution. Currently, Transformers [9, 41] have attained the state-of-the-art performance for spectral super-resolution by leveraging spectral-wise correlation through a spectral-wise self-attention mechanism. However, two types of bottlenecks exist that limit performance improvement and practical applications. First, existing Transformers predominantly focus on spectral-wise correlation while overlooking spatial-wise correlation in spectral super-resolution. Some works in other tasks [28, 15, 42, 33] attempt to model both spectral-wise and spatial-wise correlations together but often utilize separate network modules. The neglect and separation undermine the 3D nature of HSI and hinder the exploitation of unified spectral-spatial correlation. Second, existing spectral-wise self-attention mechanism always captures the full-rank correlation matrix by learning the correlation between pairs of spectral bands, i.e. tokens, in the Transformer. These characteristics result in the inability to establish linear dependence widely existing in HSI among multiple tokens.
@@
In this paper, we propose a novel Exhaustive Correlation Transformer (ECT) to model the unified spatial-spectral correlation and linear dependence, both of which we believe are crucial for spectral super-resolution. The first motivation behind our method stems from the spatial-spectral similarity in HSI. Thus, we propose a Spectral-wise Discontinuous 3D (SD3D) splitting strategy to simultaneously model unified attention along the spectral and spatial dimensions. The SD3D splitting strategy contains continuous splitting in the spatial dimension and discontinuous splitting in the spectral dimension, allowing for an effective focus on spectral-wise non-local features without disrupting the continuous structure in the spatial dimension.
The second motivation behind our methods arises from the information redundancy in HSI and its low-rank characteristic [22, 16, 55]. Thus, we propose a Dynamic Low-Rank Mapping (DLRM) module to capture the linear dependence among multiple tokens. The DLRM module simultaneously interacts among multiple tokens and maps them into a low-rank space, thereby learning a low-rank dependence map. By integrating unified spatial-spectral attention and linear dependence, our ECT can model the exhaustive correlation within HSI and achieves state-of-the-art in extensive experiments on simulated and real data.
@@
We propose a novel Exhaustive Correlation Transformer (ECT) to model unified spatial-spectral correlation and linear dependence for spectral super-resolution.
@@
We propose a Dynamic Low-Rank Mapping (DLRM) module to model the linear dependence within HSI by dynamically calculating a low-rank dependence map among multiple tokens.
The experimental results indicate that our method achieves state-of-the-art performance on both simulated and real data, with the lowest error achieved under the smallest number of parameters and the lowest inference latency.
@@ 2.1 Spectral Reconstruction
HSI acquisition is typically carried out using push-broom cameras, which is time-consuming and challenging to capture dynamic or rapidly changing scenes. To address this issue, coded aperture snapshot spectral imaging (CASSI) systems have been widely used, generating 2D measurements [51, 35], which are then processed through a series of reconstruction algorithms [48, 43, 44, 7, 10, 11, 52] to obtain HSI.
However, CASSI systems are often expensive. Reconstructing HSIs from RGB images is a cost-effective alternative. Arad et al. [3] employ sparse coding for spectral super-resolution, while Aeschbacher et al. [1] use shallow learning models and achieve improved results. Due to the presence of substantial redundant information in HSI, a low-rank prior is critical for spectral reconstruction. There are several spectral reconstruction works [56, 55, 22, 49, 16, 17, 18] inspired by the low-rank prior. Recently, Three spectral super-resolution challenges [3, 4, 5] are held and significantly inspire the research. With the development of deep learning, convolutional neural networks are widely used in the spectral super-resolution task. Shi et al. [38] propose a convolutional neural network for spectral super-resolution, which win the NTIRE 2018 Challenge on Spectral Reconstruction from RGB Images. Li et al. [26, 25] introduce the channel attention mechanism into the convolutional neural network to improve the performance. Thanks to dynamic weights and long-range correlation modeling, Cai et al. [9] are the first to introduce Transformers into the field of spectral super-resolution for modeling spectral-wise correlation and win first place in the NTIRE 2022 challenge [5]. Wang et al. [41] further improve the modeling ability of spectral-wise correlation and achieve SOTA performance recently. Though recent methods based on Transformer achieve remarkable performance improvements, they only focus on modeling spectral-wise correlation while ignoring spatial-wise correlation. Moreover, both of them neglect the critical low-rank characteristic of HSI.
@@ 3 Method
In this paper, we propose an Exhaustive Correlation Transformer (ECT), which can model unified spatial-spectral attention and linear dependence simultaneously. In this section, we first introduce the macro design of ECT. Then, we delve into the micro design within ECT.
@@ 3.2 Micro Design
As depicted in Figure 3, an Exhaustive Self-Attention Block (ESAB) comprises Exhaustive Self-Attention (ESA), Layer Normalization, and a Feed-Forward Network. The key process in ESA is summarized as follows: Firstly, the Spectral-wise Discontinuous 3D (SD3D) splitting strategy is applied to generate tokens, facilitating the exploitation of unified spatial-spectral correlation. Following that, the Low-Rank Dependence Map is generated through the Dynamic Low-Rank Mapping (DLRM) module to model linear dependence among multiple tokens. The Full-Rank Attention Map is generated through the Unified Spatial-Spectral self-Attention (USSA) module to model the independent correlation between pairs of tokens. Next, we introduce the implementation details of ESA.
@@
where denotes softmax and is a learnable parameter. , , and . normalization is performed within each Token to stabilize the training, and then the expressive power is improved by the learnable parameter . It is worth noting that the normalization and learnable parameter are designed by [2] to accommodate variable token sizes, which have been followed by abundant spectral-wise self-attention based methods [8, 9, 53, 41]. We align with these designs in this paper. Since the dimension of tokens is greater than the number of tokens in this scenario and the Softmax after the dot product, there is a lower risk of rank reduction in the attention maps, as discussed in [6]. Typically, the learned attention maps have a full rank or nearly full rank. From the optimization point of view, since the linear correlation is different in different HSIs, the loss can only be minimized when the attention mechanism learns full-rank or nearly full-rank attention maps. The experiments confirm this point as well, and we will discuss it in the supplementary material. Furthermore, the scaled dot-product attention is calculated independently between paired tokens and cannot capture the linear dependence among multiple tokens.
To address the limitation of self-attention in modeling linear dependence within HSI, we propose a Dynamic Low-Rank Mapping (DLRM) module. The details of DLRM are illustrated in Figure 4. Initially, the tokens from the multi-head () are restored in a 3D manner and subsequently spatially pooled, reducing the spatial dimensions of the tokens from to . Following this step, the tokens are flattened in a 3D fashion to form a two-dimensional matrix. Finally, interactions take place among various heads and tokens through a 1d convolution, yielding a feature () with dimensions , where is a hyperparameter and . The matrices and then undergo a Softmax function, followed by transposition and multiplication to generate a dynamic matrix, which is a low-rank matrix with a rank no greater than . The calculation of the low-rank dependence map in DLRM is expressed by
The difference between the Attention Map in self-attention and the Dependence Map in DLRM is illustrated in Figure 5. As shown in the figure, the calculation of the correlation in self-attention is token-to-token, with each row of the correlation matrix independently calculated. Each element in the matrix is solely associated with two tokens. In contrast, DLRM first facilitates information exchange among various tokens and attention heads. Therefore, each element in the Dependence Map gathers information from multiple tokens. Consequently, each element in the Dependence Map aggregates information from multiple tokens. Moreover, the Dependence Map can implicitly model the linear correlations among multiple tokens due to the low-rank nature. In summary, the Dependence Map can effectively capture the linear dependence among multiple tokens, which is absent in the Attention Map.
@@
| Method | NTIRE 2022 | NTIRE 2020 | ICVL | Params (M) | FLOPs (G) | Latency (ms) | |||
| MRAE | RMSE | MRAE | RMSE | MRAE | RMSE | ||||
| HSCNN+ [38] | 0.3814 | 0.0588 | 0.0684 | 0.0182 | 0.2322 | 0.0424 | 4.65 | 304.45 | 201 |
| HR Net [59] | 0.3476 | 0.0550 | 0.0682 | 0.0179 | 0.1139 | 0.0313 | 31.70 | 163.81 | 85 |
| AWAN [26] | 0.2191 | 0.0349 | 0.0668 | 0.0175 | 0.1040 | 0.0252 | 4.04 | 270.61 | 312 |
| Restormer [53] | 0.1833 | 0.0274 | 0.0645 | 0.0157 | 0.0945 | 0.0230 | 15.11 | 93.77 | 178 |
| DRCR [25] | 0.1823 | 0.0288 | 0.0664 | 0.0171 | 0.0763 | 0.0164 | 9.48 | 586.61 | 300 |
| MST++ [9] | 0.1645 | 0.0248 | 0.0624 | 0.0155 | 0.0691 | 0.0144 | 1.62 | 22.29 | 112 |
| HySAT [41] | 0.1599 | 0.0246 | 0.0589 | 0.0142 | 0.0654 | 0.0154 | 1.40 | 21.08 | 126 |
| ECT (Ours) | 0.1564 | 0.0236 | 0.0588 | 0.0142 | 0.0635 | 0.0142 | 1.19 | 16.75 | 82 |
@@ 4.1 Dataset
The NTIRE 2022 dataset is currently the most complex dataset for spectral super-resolution. It is captured using the Specim IQ camera and includes a wide variety of scenes and colors. The dataset contains a total of 950 available images, including 900 training images and 50 validation images. The spatial resolution of the images is , and they consist of 31 spectral channels sampled at 10 intervals, covering the wavelength range from 400 to 700.
The NTIRE 2020 dataset comprises 460 available images, including 450 training images and 10 validation images. The spatial and spectral resolutions of these images are consistent with the NTIRE 2022 dataset.
The ICVL dataset comprises 203 available HSI images. The spatial resolution is , and the spectral resolution is consistent with the NTIRE datasets. We randomly select 20 images for the validation and others for the training.
For the real data experiments, there is no available real dataset and most existing spectral super-resolution methods focus on fitting simulated data. Hence, we capture several real RGB images using FLIR Blackfly S BFS-U3-31S4 and obtained the spectral curves of the regions for validation using Specim IQ. We use HSI from the NTIRE 2022 training set and simulate RGB images to create paired training data. RGB simulation and more details about the datasets can be found in the supplementary material.
@@ 4.2 Implementation Details
For the evaluation metrics, following the NTIRE challenges, we use the Mean Relative Absolute Error (MRAE) and Root Mean Squared Error (RMSE) metrics to evaluate the performance of each network. We primarily use MRAE as the main metric and also adopt it as the training objective. RMSE is used as an auxiliary metric.
For the network training details, we utilize a batch size of 40 and employ a learning rate schedule that follows the cosine annealing scheme, decreasing from 4e4 to 1e6 over 3e5 iterations. We choose the AdamW optimizer with parameters , , and weight decay is set to 1e4. RGB images are first split into patches and undergo random rotations and flips for data augmentation before being input into the network. The codes and pretrained models will be made publicly available.